Find the measure of the following .in circle V.mAVE=90° mIVD=(2x)° and mEVS=(6X-28)°


1
Expert's answer
2021-01-25T17:12:21-0500

"V.mAVE\degree+m/VD\degree+mEVS\degree=180\degree"

"90\degree+(2X)\degree+(6X-28)\degree=180\degree"

"2X+6X=180-90+28"

"8X=62"

"X=7.75\degree"


"\therefore m/VD=2\times7.75"

"=15.5\degree"


"mEVS=(6\times7.75)-28"

"=18.5\degree"